It took three hours to row a boat 18 km against the current. The return trip with the current took 1 1/2 hours. Find the speed of the rowboat in still water.
Answer by ptaylor(1894)

Let x=boat's speed in still water and
Let y=speed of current
The boat's speed against the current (x-y) was 18km/3hr=6km/hr.
The boat's speed with the current (x+y) was 18km/1.5hr=12km/hr.
so we have:
(1) x-y=6km/hr
(2) x+y=12km/hr
adding the equations, we get:
2x=18km/hr divide both sides by 2 we have:
x=9km/hr speed of the boat in still water
substitute in (1) and we get:
9-y=6
-y=6-9
y=3km/hr speed of the current.
Hope this helps. Have a nice holiday season.----ptaylor